OpenStudy (mathhelp346):

2 1/8 + (6 2/3 ÷ 8 4/9) Please help, first person to answer correctly gets a medal :)

OpenStudy (lalaly):

\[\frac{17}{8} + (\frac{20}{3} \times \frac{9}{76})\]

OpenStudy (lalaly):

mutiply then add

OpenStudy (mathhelp346):

Ok thanks